OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 9x + 6y = 8 3x + 2y = 1

OpenStudy (anonymous):

we can solve it by eliminating x....

OpenStudy (anonymous):

by multiplying -3 to equation 2 then add the result to equation 1, we can eliminate x, thus we can solve y alone...

OpenStudy (anonymous):

can you do it @harleeekoon ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok... i tried it but it eliminate y also.... let me try a different approach....

OpenStudy (anonymous):

let us try by substitution...

jimthompson5910 (jim_thompson5910):

Orion1213 you didn't make any mistakes when you tried to eliminate x and y was eliminated as well

OpenStudy (anonymous):

yup but i'm trying to look a solution in solving y...

jimthompson5910 (jim_thompson5910):

there are no solutions if you multiply both sides of equation 2 by 3 you get 9x + 6y = 3 so we have this new system 9x + 6y = 8 9x + 6y = 3

jimthompson5910 (jim_thompson5910):

the left side is equal to some value, say z, let z = 9x + 6y therefore, we have z = 8 and z = 3 that's a contradiction. The value z can only be equal to ONE value at a time

OpenStudy (anonymous):

ok...

jimthompson5910 (jim_thompson5910):

graphing the two lines, you'll find that they are parallel lines that never intersect

OpenStudy (anonymous):

got your point... i forgot that we can graph them since they are simply a line... thanks... i hope @harleeekoon , understand the situation... i learned to you a different approach and analysis....